# DriveVLA-W0: World Models Amplify Data Scaling Law in Autonomous Driving

### v1

@@ 1 Introduction
The promise of scaling laws (Kaplan et al., 2020; Zhai et al., 2022; Baniodeh et al., 2025; Dehghani et al., 2023) presents an attractive path toward more generalized driving intelligence, with the hope that petabytes of driving data can be harnessed to train powerful foundation models. In the current landscape, two dominant paradigms exist. On one side are specialized models (Hu et al., 2022b; Jiang et al., 2023) centered around Bird’s-Eye-View (BEV) representations (Li et al., 2022; Huang et al., 2021). These models are built upon carefully designed geometric priors, which, while effective for driving-specific tasks, make it less straightforward to leverage non-driving datasets. In addition, their relatively compact architectures may constrain their potential for large-scale data scalability. In response, Vision-Language-Action (VLA) models (Fu et al., 2025; Li et al., 2025b; Zhou et al., 2025c) have emerged as a promising alternative. By leveraging large-scale Vision-Language Models (VLMs) (Wang et al., 2024; Bai et al., 2025) pretrained on internet-scale data, VLAs possess a significantly larger model size and a greater intrinsic potential for scaling.
@@
To address this supervision deficit, we propose a new training paradigm that uses world modeling (Li et al., 2024a; Wang et al., 2025; Cen et al., 2025; Chen et al., 2024) as a powerful form of self-supervision to supplement the sparse action signal. By tasking the model with predicting future images, we generate a dense and rich supervisory signal at every timestep. This objective forces the model to learn the underlying dynamics of the environment and build a rich, predictive world representation. To validate the effectiveness of our approach, we implement it across the two dominant VLA architectural families, which are primarily differentiated by their visual representation: discrete tokens versus continuous features. For VLAs that represent images as discrete visual tokens, world modeling is a natural extension. We propose an autoregressive world model to predict the sequence of discrete visual tokens of future images. For VLAs that operate on continuous features, this task is more challenging as they lack a visual vocabulary, making a direct next-token prediction approach infeasible. To bridge this gap, we introduce a diffusion world model that generates future image pixels conditioned on the vision and action features produced at the current frame.
@@ 2 Related Work
VLAs in Autonomous Driving. The development of VLMs in autonomous driving has progressed through three main stages, evolving from focusing on interpretation to integrated end-to-end VLA architectures. The first stage, language-based scene interpretation models, focused on enhancing interpretability. Models like DriveGPT4 (Xu et al., 2024) used Large Language Models (LLMs) to generate scene explanations and high-level maneuver suggestions, but without producing directly executable actions (Chen et al., 2025; Zhou et al., 2025b). The second stage, modular language-to-action frameworks, aimed to connect high-level commands with low-level control. These approaches (Zhou et al., 2025a; Yuan et al., 2024; Zhang et al., 2025; Arai et al., 2025; Gao et al., 2025) utilize multi-stage pipelines where modules are connected via non-differentiable interfaces, such as discrete textual commands, which prevent gradient back-propagation. The current frontier is End-to-End VLAs (Renz et al., 2025; 2024; Jiang et al., 2025; Hwang et al., 2024; Yang et al., 2025), which employ a unified architecture to directly map sensor inputs to trajectories. Within this paradigm, DriveMoE (Yang et al., 2025) introduced a Mixture-of-Experts (MoE) architecture, where an action decoder is serially connected after a VLM backbone. ReCogDrive (Li et al., 2025b) integrates a VLM with a diffusion planner trained via imitation and reinforcement learning to bridge the language-action gap. AutoVLA (Zhou et al., 2025c) tokenizes trajectories into action primitives, enabling a single autoregressive model to learn adaptive reasoning and planning. We follow this paradigm and propose an end-to-end VLA.
Scaling Laws in Deep Learning. Kaplan et al. (2020) were the first to systematically demonstrate that pretraining loss scales as a power law with model size, dataset size, and computational cost. Chinchilla (Hoffmann et al., 2022) then showed many LMs were undertrained and derived a compute-optimal prescription that scales model size and tokens proportionally. In computer vision, Zhai et al. (2022) charted ViT scaling law with stable training recipes, and ViT-22B (Dehghani et al., 2023) scaled ViTs to 22B parameters, verifying predictable multi-task improvements. Lin et al. (2025) conducted a large-scale study of imitation-learning data scaling in robotics, and found near–power-law gains from increasing environmental and object diversity with improved zero-shot generalization. In autonomous driving, STR (Sun et al., 2023) shows large trajectory models scale steadily in both prediction and planning, and Baniodeh et al. (2025) reported power-law improvements for joint motion forecasting and planning with large driving datasets. For end-to-end driving, Naumann et al. (2025) observe roughly log-linear gains in both open- and closed-loop metrics as training data scale increases. Zheng et al. (2024) also observe power-law improvements from data scaling in a large-scale imitation learning study. However, Zheng et al. (2024) exclusively analyzes scaling behavior within the conventional paradigm of sparse action supervision. Our work, in contrast, provides the first study of how these scaling laws are reshaped by the self-supervised signals provided by world modeling.
@@ 3.1 VLA Baseline
Input Tokenization. High-level driving language instructions ($L_{t}$) are processed using the VLM’s native tokenizer. For past actions, we use the FAST tokenizer (Pertsch et al., 2025) to convert continuous waypoint trajectories into a sequence of discrete tokens, denoted $A_{t-1}$.
VLM Backbone. At each timestep $t$, we form a deeply interleaved input sequence $S_{t}$ by concatenating multimodal chunks over a history of $H$ steps following Wang et al. (2025); Fan et al. (2025): $S_{t}=[L_{t-H},V_{t-H},A_{t-H-1},\dots,L_{t},V_{t},A_{t-1}]$. This sequence is processed autoregressively by our VLM backbone, for which we select two representative models: Emu3 (8B) (Wang et al., 2024) to handle discrete visual representations and Qwen2.5-VL (7B) (Bai et al., 2025) for continuous features, using a causal attention mask. The VLM backbone outputs the final-layer hidden states, which are then split according to their respective modalities into language ($F_{t}^{L}$), vision ($F_{t}^{V}$), and action ($F_{t}^{A}$) features.
@@
During inference, the trained model then autoregressively generates a sequence of action tokens conditioned on the context $S_{t}$. These tokens are subsequently converted back into a continuous waypoint trajectory by the FAST detokenizer (Pertsch et al., 2025).
@@ 3.2 World Modeling
Inference. While the explicit generation of visual tokens is typically bypassed during inference to ensure low latency, this capability remains valuable for visualization purposes. To generate an image, the model autoregressively samples a sequence of visual tokens, which are then passed to the MoVQGAN (Zheng et al., 2022) decoder to render the final image $\hat{I}_{t}$.
Diffusion World Model. Unlike the VQ-based counterpart, our VLA (ViT) model lacks a discrete visual vocabulary suitable for next-token prediction. We therefore introduce a Diffusion World Model, which instead provides dense supervision by training a latent diffusion model (Rombach et al., 2021) to generate future images conditioned on the VLA’s rich output features ($F_{t}^{V},F_{t}^{A}$) as Figure 2 shows. This choice to predict the future frame ($\mathbf{I}_{t+1}$) is critical: since the model is conditioned on all present features simultaneously, predicting the future is necessary to learn predictive dynamics rather than simply performing a reconstruction task.
@@ 3.3 Action Expert
MoE Architecture. (Black et al., ; Intelligence et al., 2025) While our large VLA backbone excels at representation learning, its size is prohibitive for real-time control. To address this, we introduce a lightweight action expert (500M) that operates alongside the main VLA Expert (our full VLA backbone) in a Mixture-of-Experts (MoE) architecture. The Action Expert shares a similar transformer block structure with the VLA Expert but uses a much smaller hidden dimension. This architectural similarity enables a deep and efficient fusion of information via a Joint Attention mechanism as Figure 3(a) shows. In this setup, both experts first compute their respective Query, Key, and Value matrices. These matrices are then concatenated along the token sequence dimension to create a single set of inputs for a Joint Attention operation
@@
Flow Matching Action Expert. In contrast to the discrete nature of the autoregressive approach, we also implement a continuous action generation method based on flow matching. This method learns a conditional vector field, $v_{\phi}$, that defines a direct “path” from a simple noise distribution to the complex distribution of real-world driving actions. During training, we define a simple straight-line trajectory between a random noise sample and a ground-truth action (Liu et al., 2022). The model is then optimized via a mean squared error loss to predict a vector field $v_{\phi}$ that aligns with this trajectory at each point, conditioned on the multimodal context $\mathbf{c}_{t}$. For inference, we simply start with a new noise sample and follow the learned vector field for a fixed number of steps using a numerical ODE solver. This process deterministically transforms the noise into a precise, continuous action that lies on the learned data manifold.
@@ 4 Experiment
Table 1: Comparison with state-of-the-art methods on the NAVSIM v1. NC: no at-fault collision. DAC: drivable area compliance. TTC: time-to-collision. C.: comfort. EP: ego progress. PDMS: the predictive driver model score. Abbreviations: 1x Cam (single front-view camera), Nx Cam (surround-view cameras), L (LiDAR). : Using the query-based action expert with multiple trajectory anchors following Li et al. (2024b). †: Using the AR action expert with the best-of-N (N=6) strategy following Zhou et al. (2025c).
@@
UniAD (Hu et al., 2022b) CVPR’23 6x Cam 97.8 91.9 92.9 100.0 78.8 83.4
TransFuser (Prakash et al., 2021) TPAMI’23 3x Cam + L 97.7 92.8 92.8 100.0 79.2 84.0
PARA-Drive (Weng et al., 2024) CVPR’24 6x Cam 97.9 92.4 93.0 99.8 79.3 84.0
LAW (Li et al., 2024a) ICLR’25 1x Cam 96.4 95.4 88.7 99.9 81.7 84.6
Hydra-MDP (Li et al., 2024b) arXiv’24 3x Cam + L 98.3 96.0 94.6 100.0 78.7 86.5
DiffusionDrive (Liao et al., 2025) CVPR’25 3x Cam + L 98.2 96.2 94.7 100.0 82.2 88.1
WoTE (Li et al., 2025a) ICCV’25 3x Cam + L 98.5 96.8 94.4 99.9 81.9 88.3
@@
AutoVLA (Zhou et al., 2025c) NeurIPS’25 3x Cam 98.4 95.6 98.0 99.9 81.9 89.1
ReCogDrive (Li et al., 2025b) arXiv’25 3x Cam 98.2 97.8 95.2 99.8 83.5 89.6
DriveVLA-W0 - 1x Cam 98.7 99.1 95.3 99.3 83.3 90.2
AutoVLA† (Zhou et al., 2025c) NeurIPS’25 3x Cam 99.1 97.1 97.1 100.0 87.6 92.1
DriveVLA-W0† - 1x Cam 99.3 97.4 97.0 99.9 88.3 93.0
@@
TransFuser (Prakash et al., 2021) 96.9 89.9 97.8 99.7 87.1 95.4 92.7 98.3 87.2 76.7
HydraMDP++ (Li et al., 2024b) 97.2 97.5 99.4 99.6 83.1 96.5 94.4 98.2 70.9 81.4
DriveSuprem (Yao et al., 2025) 97.5 96.5 99.4 99.6 88.4 96.6 95.5 98.3 77.0 83.1
ARTEMIS (Feng et al., 2025) 98.3 95.1 98.6 99.8 81.5 97.4 96.5 98.3 - 83.1
DiffusionDrive (Liao et al., 2025) 98.2 95.9 99.4 99.8 87.5 97.3 96.8 98.3 87.7 84.5
@@ 4.1 Datasets
NAVSIM. We use the NAVSIM (Dauner et al., 2024) benchmark derived from OpenScene (Contributors, 2023), for evaluating performance in safety-critical scenarios.
@@
NAVSIM v2 (Cao et al., 2025) includes several components, categorized as penalties or weighted subscores. Key metrics are No at-fault Collision (NC), Drivable Area Compliance (DAC), Driving Direction Compliance (DDC), Traffic Light Compliance (TLC), Ego Progress (EP), Time to Collision (TTC), Lane Keeping (LK), History Comfort (HC), and Extended Comfort (EC). NAVSIM v2 uses the Extended Predictive Driver Model Score (EPDMS) to evaluate model performance: $\text{EPDMS}=\text{NC}\times\text{DAC}\times\text{DDC}\times\text{TLC}\times\frac{5\times\text{EP}+5\times\text{TTC}+2\times\text{LK}+2\times\text{HC}+2\times\text{EC}}{16}$.
@@ 4.2 Implementation Details
NAVSIM. For experiments on the NAVSIM benchmark, models are pretrained on NuPlan (Caesar et al., 2021) for 8k steps and then fine-tuned on NAVSIM for 4k steps, processing 256x144 images. The training is conducted on 8 NVIDIA L20 GPUs with a global batch size of 48. We used the AdamW optimizer with a cosine learning rate schedule, an initial learning rate of 2e42e-4, and bfloat16 mixed-precision. For our ablation studies, we select DriveVLA-W0 (VQ) as the default model due to its architectural simplicity.
@@
Re-implemented TransFuser. To provide a solid baseline for our in-house dataset, we adapt and re-implement the well-known TransFuser (Prakash et al., 2021) architecture. For a fair comparison with our single-camera setup, we modify Latent-TransFuser (an image-only variant of TransFuser) to process a single front-view image instead of its original multi-view input. To investigate the impact of model size, we implement two versions: a 50M-parameter model and a 7B-parameter model. The smaller TransFuser-50M uses a standard ResNet-34 backbone. The larger TransFuser-7B employs a ViT-7B backbone, initialized with pretrained weights from DINOv3 (Siméoni et al., 2025).

### v2

@@ 1 Introduction
The promise of scaling laws (kaplan2020scaling; zhai2022scaling; baniodeh2025scaling; dehghani2023scaling) presents an attractive path toward more generalized driving intelligence, with the hope that petabytes of driving data can be harnessed to train powerful foundation models. In the current landscape, two dominant paradigms exist. On one side are specialized models (hu2022uniad; jiang2023vad) centered around Bird’s-Eye-View (BEV) representations (li2022bevformer; huang2021bevdet). These models are built upon carefully designed geometric priors, which, while effective for driving-specific tasks, make it less straightforward to leverage non-driving datasets. In addition, their relatively compact architectures may constrain their potential for large-scale data scalability. In response, Vision-Language-Action (VLA) models (fu2025orion; li2025recogdrive; zhou2025autovla) have emerged as a promising alternative. By leveraging large-scale Vision-Language Models (VLMs) (wang2024emu3; Qwen2.5-VL) pretrained on internet-scale data, VLAs possess a significantly larger model size and a greater intrinsic potential for scaling.
@@
To address this supervision deficit, we harness the power of world modeling (law; univla; worldvla; chen2024drivinggpt), integrating it as a dense self-supervised objective to supplement the sparse action signal. By tasking the model with predicting future images, we generate a dense and rich supervisory signal at every timestep. This objective forces the model to learn the underlying dynamics of the environment and build a rich, predictive world representation. To validate the effectiveness of our approach, we implement it across the two dominant VLA architectural families, which are primarily differentiated by their visual representation: discrete tokens versus continuous features. For VLAs that represent images as discrete visual tokens, world modeling is a natural extension. We propose an autoregressive world model to predict the sequence of discrete visual tokens of future images. For VLAs that operate on continuous features, this task is more challenging as they lack a visual vocabulary, making a direct next-token prediction approach infeasible. To bridge this gap, we introduce a diffusion world model that generates future image pixels conditioned on the vision and action features produced at the current frame.
@@ 2 Related Work
VLAs in Autonomous Driving. The development of VLMs in autonomous driving has progressed through three main stages, evolving from focusing on interpretation to integrated end-to-end VLA architectures. The first stage, language-based scene interpretation models, focused on enhancing interpretability. Models like DriveGPT4 (drivegpt4) used Large Language Models (LLMs) to generate scene explanations and high-level maneuver suggestions, but without producing directly executable actions (ts-vlm; dynrsl-vlm). The second stage, modular language-to-action frameworks, aimed to connect high-level commands with low-level control. These approaches (opendrivevla; rag-drive; safeauto; covla; langcoop) utilize multi-stage pipelines where modules are connected via non-differentiable interfaces, such as discrete textual commands, which prevent gradient back-propagation. The current frontier is End-to-End VLAs (simlingo; carllava; diffvla; emma; drivemoe; zeng2025futuresightdrive), which employ a unified architecture to directly map sensor inputs to trajectories. Within this paradigm, DriveMoE (drivemoe) introduced a Mixture-of-Experts (MoE) architecture, where an action decoder is serially connected after a VLM backbone. ReCogDrive (li2025recogdrive) integrates a VLM with a diffusion planner trained via imitation and reinforcement learning to bridge the language-action gap. AutoVLA (zhou2025autovla) tokenizes trajectories into action primitives, enabling a single autoregressive model to learn adaptive reasoning and planning. We follow this paradigm and propose an end-to-end VLA.
World Models in Driving and Robotics Research on world models typically follows two distinct philosophies: utilizing them as data synthesizers for simulation, or employing them as an auxiliary objective for representation learning. The first stream focuses on generation, aiming to synthesize high-quality driving data. Prominent works like GAIA-1 (hu2023gaia), DrivingGPT (chen2024drivinggpt), and Doe-1 (zheng2024doe) generate future images or multi-modal tokens to create realistic scenarios, while Copilot4D (zhang2023copilot4d) specifically predicts future discrete visual tokens. Conversely, the second stream leverages world modeling as a self-supervised objective to enhance representation learning. While VaVAM (bartoccioni2025vavim) demonstrates the value of generative video pretraining, it lacks action conditioning. Addressing this, UniVLA (univla) and WorldVLA (worldvla) incorporate actions into autoregressive prediction to learn robust policies. In autonomous driving, LAW (law) pioneers this paradigm using a latent world model. Distinct from LAW’s latent prediction, our work supervises the model to predict future images, providing a richer and more direct dense learning signal for the VLA backbone.
@@ 3.1 VLA Baseline
Input Tokenization. High-level driving language instructions ($L_{t}$) are processed using the VLM’s native tokenizer. For past actions, we use the FAST tokenizer (pifast) to convert continuous waypoint trajectories into a sequence of discrete tokens, denoted $A_{t-1}$.
VLM Backbone. At each timestep $t$, we form a deeply interleaved input sequence $S_{t}$ by concatenating multimodal chunks over a history of $H$ steps following univla; interleavevla: $S_{t}=[L_{t-H},V_{t-H},A_{t-H-1},\dots,L_{t},V_{t},A_{t-1}]$. This sequence is processed autoregressively by our VLM backbone, for which we select two representative models: Emu3 (8B) (wang2024emu3) to handle discrete visual representations and Qwen2.5-VL (7B) (Qwen2.5-VL) for continuous features, using a causal attention mask. The VLM backbone outputs the final-layer hidden states, which are then split according to their respective modalities into language ($F_{t}^{L}$), vision ($F_{t}^{V}$), and action ($F_{t}^{A}$) features.
@@
During inference, the trained model then autoregressively generates a sequence of action tokens conditioned on the context $S_{t}$. These tokens are subsequently converted back into a continuous waypoint trajectory by the FAST detokenizer (pifast).
@@ 3.2 World Modeling
Inference. While the explicit generation of visual tokens is typically bypassed during inference to ensure low latency, this capability remains valuable for visualization purposes. To generate an image, the model autoregressively samples a sequence of visual tokens, which are then passed to the MoVQGAN (movqgan) decoder to render the final image $\hat{I}_{t}$.
Diffusion World Model. Unlike the VQ-based counterpart, our VLA (ViT) model lacks a discrete visual vocabulary suitable for next-token prediction. We therefore introduce a Diffusion World Model, which instead provides dense supervision by training a latent diffusion model (stablediffusion; ross) to generate future images conditioned on the VLA’s rich output features ($F_{t}^{V},F_{t}^{A}$) as Figure 2 shows. This choice to predict the future frame ($\mathbf{I}_{t+1}$) is critical: since the model is conditioned on all present features simultaneously, predicting the future is necessary to learn predictive dynamics rather than simply performing a reconstruction task.
@@ 3.3 Action Expert
MoE Architecture. (pi0; pi05) While our large VLA backbone excels at representation learning, its size is prohibitive for real-time control. To address this, we introduce a lightweight action expert (500M) that operates alongside the main VLA Expert (our full VLA backbone) in a Mixture-of-Experts (MoE) architecture. The Action Expert shares a similar transformer block structure with the VLA Expert but uses a much smaller hidden dimension. This architectural similarity enables a deep and efficient fusion of information via a Joint Attention mechanism as Figure 3(a) shows. In this setup, both experts first compute their respective Query, Key, and Value matrices. These matrices are then concatenated along the token sequence dimension to create a single set of inputs for a Joint Attention operation
@@
Flow Matching Action Expert. In contrast to the discrete nature of the autoregressive approach, we also implement a continuous action generation method based on flow matching. This method learns a conditional vector field, $v_{\phi}$, that defines a direct “path” from a simple noise distribution to the complex distribution of real-world driving actions. During training, we define a simple straight-line trajectory between a random noise sample and a ground-truth action (liu2022flow). The model is then optimized via a mean squared error loss to predict a vector field $v_{\phi}$ that aligns with this trajectory at each point, conditioned on the multimodal context $\mathbf{c}_{t}$. For inference, we simply start with a new noise sample and follow the learned vector field for a fixed number of steps using a numerical ODE solver. This process deterministically transforms the noise into a precise, continuous action that lies on the learned data manifold.
@@ 4 Experiment
Table 1: Comparison with state-of-the-art methods on the NAVSIM v1. NC: no at-fault collision. DAC: drivable area compliance. TTC: time-to-collision. C.: comfort. EP: ego progress. PDMS: the predictive driver model score. Abbreviations: 1x Cam (single front-view camera), Nx Cam (surround-view cameras), L (LiDAR). : Using the query-based action expert. †: Using the query-based action expert with multiple trajectory anchors following hydra. ‡: Using the AR action expert with the best-of-N (N=6) strategy following zhou2025autovla.
@@
UniAD (hu2022uniad) CVPR’23 6x Cam 97.8 91.9 92.9 100.0 78.8 83.4
TransFuser (transfuser) TPAMI’23 3x Cam + L 97.7 92.8 92.8 100.0 79.2 84.0
PARA-Drive (paradrive) CVPR’24 6x Cam 97.9 92.4 93.0 99.8 79.3 84.0
LAW (law) ICLR’25 1x Cam 96.4 95.4 88.7 99.9 81.7 84.6
Hydra-MDP (hydra) arXiv’24 3x Cam + L 98.3 96.0 94.6 100.0 78.7 86.5
DiffusionDrive (diffdrive) CVPR’25 3x Cam + L 98.2 96.2 94.7 100.0 82.2 88.1
WoTE (wote) ICCV’25 3x Cam + L 98.5 96.8 94.4 99.9 81.9 88.3
@@
DriveVLA-W0 - 1x Cam 98.7 96.2 95.5 100.0 82.2 88.4
AutoVLA (zhou2025autovla) NeurIPS’25 3x Cam 98.4 95.6 98.0 99.9 81.9 89.1
ReCogDrive (li2025recogdrive) arXiv’25 3x Cam 98.2 97.8 95.2 99.8 83.5 89.6
[…]
AutoVLA† (zhou2025autovla) NeurIPS’25 3x Cam 99.1 97.1 97.1 100.0 87.6 92.1
DriveVLA-W0‡ - 1x Cam 99.3 97.4 97.0 99.9 88.3 93.0
@@
TransFuser (transfuser) 96.9 89.9 97.8 99.7 87.1 95.4 92.7 98.3 87.2 76.7
HydraMDP++ (hydra) 97.2 97.5 99.4 99.6 83.1 96.5 94.4 98.2 70.9 81.4
DriveSuprem (yao2025drivesuprim) 97.5 96.5 99.4 99.6 88.4 96.6 95.5 98.3 77.0 83.1
ARTEMIS (feng2025artemis) 98.3 95.1 98.6 99.8 81.5 97.4 96.5 98.3 - 83.1
DiffusionDrive (diffdrive) 98.2 95.9 99.4 99.8 87.5 97.3 96.8 98.3 87.7 84.5
@@ 4.1 Datasets
NAVSIM. We use the NAVSIM (navsim) benchmark derived from OpenScene (openscene), for evaluating performance in safety-critical scenarios.
@@
NAVSIM v2 (cao2025navsimv2) includes several components, categorized as penalties or weighted subscores. Key metrics are No at-fault Collision (NC), Drivable Area Compliance (DAC), Driving Direction Compliance (DDC), Traffic Light Compliance (TLC), Ego Progress (EP), Time to Collision (TTC), Lane Keeping (LK), History Comfort (HC), and Extended Comfort (EC). NAVSIM v2 uses the Extended Predictive Driver Model Score (EPDMS) to evaluate model performance: $\text{EPDMS}=\text{NC}\times\text{DAC}\times\text{DDC}\times\text{TLC}\times\frac{5\times\text{EP}+5\times\text{TTC}+2\times\text{LK}+2\times\text{HC}+2\times\text{EC}}{16}$.
@@ 4.2 Implementation Details
NAVSIM. For experiments on the NAVSIM benchmark, models are pretrained on NuPlan (nuplan) for 8k steps and then fine-tuned on NAVSIM for 4k steps, processing 256x144 images. The training is conducted on 8 NVIDIA L20 GPUs with a global batch size of 48. We used the AdamW optimizer with a cosine learning rate schedule, an initial learning rate of 2e42e-4, and bfloat16 mixed-precision. For our ablation studies, we select DriveVLA-W0 (VQ) as the default model due to its architectural simplicity.
@@
Re-implemented TransFuser. To provide a solid baseline for our in-house dataset, we adapt and re-implement the well-known TransFuser (transfuser) architecture. For a fair comparison with our single-camera setup, we modify Latent-TransFuser (an image-only variant of TransFuser) to process a single front-view image instead of its original multi-view input. To investigate the impact of model size, we implement two versions: a 50M-parameter model and a 7B-parameter model. The smaller TransFuser-50M uses a standard ResNet-34 backbone. The larger TransFuser-7B employs a ViT-7B backbone, initialized with pretrained weights from DINOv3 (simeoni2025dinov3).